### v2

# Multi-Pair Temporal Sentence Grounding via Multi-Thread Knowledge Transfer Network

Abstract
Given some video-query pairs with untrimmed videos and sentence queries, temporal sentence grounding (TSG) aims to locate query-relevant segments in these videos. Although previous respectable TSG methods have achieved remarkable success, they train each video-query pair separately and ignore the relationship between different pairs. We observe that the similar video/query content not only helps the TSG model better understand and generalize the cross-modal representation, but also assists the model in locating some complex video-query pairs. Previous methods follow a single-thread framework that cannot co-train different pairs and usually spends much time re-obtaining redundant knowledge, limiting their real-world applications. To this end, in this paper, we pose a brand-new setting: Multi-Pair TSG, which aims to co-train these pairs. In particular, we propose a novel video-query co-training approach, Multi-Thread Knowledge Transfer Network, to locate a variety of video-query pairs effectively and efficiently. Firstly, we mine the spatial and temporal semantics across different queries to cooperate with each other. To learn intra- and inter-modal representations simultaneously, we design a cross-modal contrast module to explore the semantic consistency by a self-supervised strategy. To fully align visual and textual representations between different pairs, we design a prototype alignment strategy to 1) match object prototypes and phrase prototypes for spatial alignment, and 2) align activity prototypes and sentence prototypes for temporal alignment. Finally, we develop an adaptive negative selection module to adaptively generate a threshold for cross-modal matching. Extensive experiments show the effectiveness and efficiency of our proposed method.
Introduction
Temporal sentence grounding (TSG) (Gao et al. 2017; Fang et al. 2022, 2023b, 2023c, 2024b, 2023a, 2024c; Li et al. 2023, 2024; Yu et al. 2024c; Ning et al. 2024, 2023a, 2023b; Wang et al. 2025a) is an important yet challenging multi-modal task, which has received increasing attention in recent years due to its wide potential applications, such as video understanding (Liu et al. 2024c, 2023e, d, b; Wang et al. 2024c; Hu et al. 2023a; Fei et al. 2024d, c, b, a; Wu et al. 2024; Wang et al. 2024b, 2023, 2022a; Zhang et al. 2024a, b; Yu et al. 2022, 2025) and human-computer interaction (Liu et al. 2024f, g, 2023g, 2023c, 2023d; Tang et al. 2022b, a, 2023a, 2024a; Feng et al. 2023b, a; Zhao et al. 2024a, b; Wang et al. 2024a; Yu et al. 2024b, 2023). By complex multi-modal interactions and complicated context information, TSG targets the challenging problem of locating a variety of sentence queries about a video, which requires the designed models to understand both natural language and long video, including reasoning about activities, objects, sequence of events, and interactions within the video (Hu et al. 2022; Wei et al. 2023; Zhao et al. 2021, 2017, 2022, 2018a, 2018b; Jia et al. 2019, 2022b, 2022c, 2022a, 2024a, 2021, 2024d, 2020, 2023, 2024b, 2024e, 2024c; Gao et al. 2021, 2022, 2024a, 2024b, 2024c; Fang et al. 2025a; Hu et al. 2020b, 2021b, 2023b, 2023b; Hu, Liu, and Wu 2021; Hu et al. 2020c, 2024d, a, 2024d, 2023d, 2024e, 2024c, 2021a, 2024b, 2024e; Hu and Su 2024; Hu et al. 2024e, 2023f, 2023e, 2023c; Li et al. 2022, 2025). As shown in Figure 1(a), given an untrimmed video and a sentence query, TSG aims to determine the segment boundaries that contain the query-relevant activity (Qu et al. 2020; Dong et al. 2022b; Sun et al. 2024; Qu et al. 2024a; Zheng et al. 2023; Qu et al. 2024b).
Figure 1: (a) Example of temporal sentence grounding (TSG). (b) Previous TSG models regard each video-query pair independently. (c) Our proposed model explores the semantic relationship between different pairs to reduce the modality gap.
Most previous TSG works (Xiu et al. 2024; Liu et al. 2022, 2020, 2021; Ji et al. 2023c, b, 2024d, 2024b, 2024a, a) refer to a fully-supervised setting, where each frame is manually labeled as query-relevant or not. To avoid using such labour-intensive frame-level annotations, some recent works (Fang et al. 2025b, 2020; Fang and Hu 2020; Fang et al. 2021b, a; Fang, Easwaran, and Genest 2024; Fang et al. 2024a; Liu et al. 2024j, 2023f, h, i; Zhang et al. 2024c; Yu et al. 2024a; Zhang et al. 2024d, 2023d, 2023c, 2023b, 2022; Guo et al. 2024; Dong et al. 2022a) explore a weakly-supervised setting with only the video-query correspondence to alleviate the reliance to a certain extent. Despite the remarkable performance, fully- and weakly-supervised methods only treat each video-query pair independently and ignore the semantic relationship between different video-query pairs, as shown in Figure 1(b). Since a video often corresponds to multiple queries, their treatment will repeatedly extract video features and repetitively conduct complex multi-modal calculations, resulting in weak efficiency. Besides, ignoring the semantic relationship between different video-query pairs might miss important spatio-temporal semantic consistency (e.g., common noun/appearance “woman” and temporal relationship “continues to” in Figure 2), which limits their effectiveness. Thus, an effective and efficient model is expected to explore the latent semantics relationship between different video-query pairs.
Hence, we pose a novel task: can we co-train multiple video-query pairs and transfer the grounding knowledge from a pair to another pair? We demonstrate this brand-new task “multi-pair TSG” (MP-TSG) in Figure 1(c). To the best of our knowledge, there is no such setting proposed in existing works. To address this brand-new and challenging setting, we propose a novel multi-thread framework to co-train different pairs. We notice that the semantic relationship between different video-query pairs includes four aspects: query-to-query relationship, video-to-query relationship, cross-modal spatial relationship (object-to-phrase) and cross-modal temporal relationship (activity-to-sentence). Specifically, we first mine the shared spatial semantics and temporal relationships across different sentence queries to assist with each other in the TSG task. To mine the intra-modal information and obtain inter-modal representation simultaneously, we then design a cross-modal contrast module to explore the global-level semantic consistency between videos and queries by a self-supervised strategy. Moreover, we design an adaptive negative selection module to adaptively generate a dynamic threshold for cross-modal matching. To sufficiently align fine-grained visual information and fine-grained textual information from spatial and temporal perspectives, we design a prototype alignment strategy to 1) match the object prototypes and phrase prototypes to align appearance representations across modalities, and 2) align activity prototypes and sentence prototypes to integrate motion representations between different modalities. Our main contributions are as follows:
  • We pose and address a brand-new task: MP-TSG, which aims to co-train multiple video-query pairs by exploring the semantic relationships between different pairs to assist with each other. We propose a novel multi-thread framework to co-train different pairs by mining the relationships in four aspects: query-to-query, video-to-query, object-to-phrase, activity-to-sentence.
  • We propose a novel cross-modal prototype alignment module to explore the semantic relationship between different queries/videos. Besides, to deeply explore cross-modal matching, we design an adaptive negative selection module to automatically generate a dynamic threshold for semantically matching video-query pairs.
  • Extensive experiments on three challenging benchmarks demonstrate both effectiveness and efficiency of our method. More importantly, our method can serve as a plug-and-play module for state-of-the-art methods to enhance their effectiveness and efficiency.
Related Works
Fully-supervised TSG. Temporal sentence grounding (TSG) (Anne Hendricks et al. 2017; Gao et al. 2017; Ju et al. 2024, 2020, 2023; Liang et al. 2024a, b, 2023) aims at locating the most relevant segments from long videos corresponding to the given sentence descriptions. Traditional TSG methods (Gao et al. 2017; Xiong et al. 2024; Rao et al. 2021b, a; Jiang et al. 2023; Zhang et al. 2023a; Wang et al. 2025c, b; Tang et al. 2024d, 2022c, 2021, 2022d, 2023b, b, c) typically utilize a propose-and-rank approach to make predictions based on interacted multi-modal features. Some proposal-free methods (Zhang et al. 2020a; Lin et al. 2024b, a, c, 2025; Wang et al. 2021a, b, 2020) are proposed to directly predict the temporal locations of the target segment without generating proposals.
Weakly-supervised TSG. The above fully-supervised methods heavily rely on the datasets that require numerous manually labeled annotations for training. To ease human labeling efforts, recent works (Mithun, Paul, and Roy-Chowdhury 2019; Liu et al. 2023a, b, 2024a; Wen et al. 2023) consider a weakly-supervised setting to only access the information of matched video-query pairs without accurate segment boundaries. However, their performance is less satisfactory with such weak supervision.
Many semantic relationships between different queries/videos are not explored in previous methods, leading to repeated training and much computational cost. Unlike them, we introduce a brand-new setting, MP-TSG, where different queries and videos can be co-trained to reduce the cross-modal gap between video and query.
Methodology
Figure 2: Overview of our proposed MKTN for the MP-TSG task. Given some video-query pairs (e.g., the first and second videos correspond to one and two queries respectively), we first utilize video and query encoders to extract corresponding features. Then, we feed these features into four carefully-designed modules to fully explore the semantic relationships between videos and queries. In the cross-sentence semantic mining module, we mine the query-to-query relationship based on the cross-modal memory. For the adaptive video-query matching module, we adaptively learn the cross-modal semantic consistency with video-to-query relationship by a dynamic threshold $\phi$ and a contrastive loss $\mathcal{L}_{CL}$. In the object-phrase prototype matching module, we align appearance representations across modalities based on appearance and phrase prototypes. Similarly, we integrate motion representations by aligning activity and sentence prototypes. Best viewed in color.
Problem definition. Given $M_{q}$ video-query pairs $\{V_{p},Q_{p}\}_{p=1}^{M_{q}}$, previous TSG methods aim to localize the query-described activity segment in the video for each video-query pair, where $V_{p}$ and $Q_{p}$ are the corresponding video and query, respectively. They independently regard each video-query pair with ignoring the semantic relationship between different queries and videos, and repeat the grounding process $M_{q}$ times. Different from them, we pose a practical yet challenging setting, Multi-Pair TSG (MP-TSG), which aims to co-train multiple video-query pairs for effective and efficient grounding.
Pipeline. To tackle the MP-TSG task, we propose a novel framework in Figure 2. The semantic relationship between different video-query pairs includes four aspects: query-to-query relationship, video-to-query relationship, cross-modal spatial relationship (object-to-phrase) and cross-modal temporal relationship (activity-to-sentence). The first two relationships are global-level, and we can determine whether any video and any query are related or not. The last two relationships are local-level, which aligns the cross-modal semantics from spatial and temporal perspectives respectively for precise video grounding.
Preparation
Video encoder. Given $M_{v}$ videos $\{V_{1},\cdots,V_{M_{v}}\}$, we first follow previous work (Gao et al. 2017) to extract its frame-wise features by a pre-trained 3D-CNN network (Tran et al. 2015), and then employ a multi-head self-attention (Vaswani et al. 2017) module to capture the long-range dependencies among video frames. For the $a$-th video with $N_{v}$ frames, we denote the extracted video features as ${V}_{a}=\{{v}_{g}^{i},{v}_{a_{1}}^{i},\cdots,{v}_{a_{C}}^{i}\}_{i=1}^{N_{v}}\in\mathbb{R}^{N_{v}\times(C+1)\times d}$, where $d$ is the feature dimension, $C$ is the patch number, ${v}_{g}^{i}$ is the global feature of the $i$-th frame.
Text encoder. Similarly, given $M_{q}$ queries $\{Q_{1},\dots,Q_{M_{q}}\}$, by feeding any query $Q_{j}$ to the pretained Glove network (Pennington, Socher, and Manning 2014), we can obtain the word-level features ${Q}_{j}=\{{q}_{1}^{j},\cdots,{q}_{N_{q}}^{j}\}\in\mathbb{R}^{N_{q}\times d}$, where $N_{q}$ is the word number. To extract the semantic of the whole sentence, the Skip-thought parser (Kiros et al. 2015) is employed to capture the query-level feature ${q}_{e}^{j}\in\mathbb{R}^{d}$.
Cross-Sentence Semantic Mining
Although previous TSG works (Anne Hendricks et al. 2017; Gao et al. 2017) try to fully understand textual information and visual information (Zhao et al. 2024a; Xiong et al. 2024; Cai et al. 2025; Liu et al. 2024e; Tang et al. 2025, 2024d; Lei et al. 2025; Yang et al. 2025; Liu et al. 2024g; Zhang et al. 2025a, b; Hu et al. 2024a, 2025), they often ignore the semantic relationship between different sentences. The semantic relationship includes 1) the temporal information between different segments in the same video, and 2) the contextual information among different sentences. To sufficiently mine these query-to-query relationships, we extract query-level features $F_{q}=\{q_{e}^{j}\}_{j=1}^{M_{q}}$ from the multiple queries rather than learnable embeddings in previous works. Thus, we aim to model the query-level contexts, and explore the temporally and contextually related queries of each query. For example, in Figure 2, “The woman starts doing some ballet moves.” and “She continues to dance to ballet moves as the little girl walks away.” share the same semantics (“woman” and “ballet moves”), and contains the temporal relationship (“start” and “continues to”). The shared semantics and temporal relationship will assist the grounding task of each sentence.
Given the query-level features $F_{q}$, we first encode temporal information by the position embedding layer, and then conduct the interactions among queries by the self-attention layers. After that, we extract a textual feature for each query to represent corresponding events from the multi-modal memory by cross-attention layers. Considering that some sentences might share the same particular words (e.g., “woman” in Figure 2), we extract the hierarchical textual features to conduct the cross-granularity interactions, which makes the decoder learn more contextual information. Finally, we compute the timestamps of each query-wise feature by a parallel regression layer. The above procedures are formulated as $T=MLP(\tau(F_{q},F_{mem}))$, where $T=\{(t_{s}^{j},t_{e}^{j})\}_{j=1}^{M_{q}}$ denotes the ground-truth start and end timestamps ($t_{s}^{j},t_{e}^{j}$) for $M_{q}$ queries; $\tau$ denotes the decoder of the transformer to conduct the query-level position embedding; $F_{mem}=[F_{v}^{j};F_{w}^{j}]_{j=1}^{M_{q}}$ denotes the multi-modal memory, where $F_{v}^{j}=\{{v}_{a_{1}}^{j},\cdots,{v}_{a_{C}}^{j}\}$ is the frame-level feature of the video paired with query $Q_{j}$, $F_{w}^{j}=\{{q}_{1}^{j},\cdots,{q}_{N_{q}}^{j}\}$ is the word-level feature, [;][$\cdot$;$\cdot$][ ⋅ ; ⋅ ] denotes the concatenating operation.
Adaptive Video-Query Matching
Adaptive negative selection. Common TSG datasets inherently treat a set of video-query pairs as positive matches. Negatives are assigned under the assumption that all non-corresponding pairs are semantically distinct. However, some queries labeled as negative may indeed partially or accurately align with a video, constituting false negatives. Therefore, we design a dynamic threshold-based negative selection strategy to adaptively select negatives. For a given query $Q_{j}$, the negative videos are selected as:
$\mathcal{N}_{q}=\{V_{i}|S_{ij}<\phi\}\cap\mathcal{N},{S}_{ij}={V_{i}}({Q_{j}}{W}_{S})^{\top}\in\mathbb{R}^{N_{v}\times N_{q}},$ (1)
where $S_{ij}$ denotes the similarity between video $i$ and query $j$; ${W}_{S}\in\mathbb{R}^{d\times d}$ projects the query features into the same latent space as the video; $\mathcal{N}$ is the original negative set; $\phi$ is a dynamic threshold:
$\phi=\phi_{final}-(\phi_{final}-\phi_{intial})\cos(r\pi+1),$ (2)
where $\phi_{intial}$ and $\phi_{final}$ are the thresholds at the start and end of training respectively, $r\in[0,1]$ is the percentage of the training process. The design of the cosine annealing threshold where $\phi_{final}>\phi_{intial}$ is based on the intuition that the model has a higher confidence level in later training. In the early stages of training, negatives with relatively high similarity scores can be reliably regarded as false negatives.
Self-weighted cross-modal contrast. To mine the intra-modal information and obtain inter-modal representations simultaneously, we design a cross-modal contrast module to explore semantic consistency. As shown in Figure 2, we map the word-level textual feature $F_{w}$ and frame-level feature $F_{v}$ into a shared subspace for semantic alignment. Especially, we introduce the transformer encoder $\varsigma$ to generate the transferred word-level textual feature $F^{\prime}_{w}$ and video feature $F^{\prime}_{v}$ as: $F^{\prime}_{w}=Norm(\varsigma(F_{w})),F^{\prime}_{v}=Norm(\varsigma(F_{v}))$.
We construct a triplet tuple $(F^{\prime+}_{v},F^{\prime+}_{w},F^{\prime-}_{w})$ to denote the pair relationship across queries, where $(F^{\prime+}_{v},F^{\prime+}_{w})$ is a positive pair and $(F^{\prime+}_{v},F^{\prime-}_{w})$ is a negative pair. To pull the positive pairs $(F^{\prime+}_{v},F^{\prime+}_{w})$ together and push the negative pairs $(F^{\prime+}_{v},F^{\prime-}_{w})$ away, we map the word-level textual feature $F_{w}$ and frame-level feature $F_{v}$ into a shared subspace for semantic alignment. Similarly, we can obtain the inter-video relationship: $(F^{\prime+}_{w},F^{\prime+}_{v},F^{\prime-}_{v})$, where $(F^{\prime+}_{w},F^{\prime+}_{v})$ is a positive pair and $(F^{\prime+}_{w},F^{\prime-}_{v})$ is a negative pair. Thus, we design the following contrastive loss for self-supervision:
$\mathcal{L}_{CL}\!=\!\sum\nolimits_{F^{\prime+}_{v},F^{\prime+}_{w}}\!\{\theta\!\sum\nolimits_{F^{\prime-}_{w}}\!\max[0,\!\phi\!-\!S_{(F^{\prime+}_{v},F^{\prime+}_{w})}\!+\!S_{(F^{\prime+}_{v},F^{\prime-}_{w})}]$
$+(1-\theta)\sum\nolimits_{F^{\prime-}_{v}}\max[0,\phi-S_{(F^{\prime+}_{w},F^{\prime+}_{v})}+S_{(F^{\prime+}_{w},F^{\prime-}_{v})}]\},$ (3)
where $\phi$ is the dynamic threshold in Eq. (2); the scoring function $S_{(\cdot,\cdot)}$ measures the similarity between the visual feature and textual feature in the joint space; $\theta\in(0,1)$ is a regularized parameter to balance the significance between negative videos and negative queries. We can obtain $\theta$ by:
$\frac{\theta}{1-\theta}=\frac{(||F^{\prime+}_{v}-F^{\prime-}_{w}||_{2}-||F^{\prime+}_{v}-F^{\prime+}_{w}||_{2})^{2}}{(||F^{\prime+}_{v}-F^{\prime-}_{w}||_{2}-||F^{\prime+}_{v}-F^{\prime+}_{w}||_{2})^{2}}.$ (4)
Object-Phrase Prototype Matching
The query subjectively describes a video activity, resulting in partially matched video-query pairs. Although a single patch-activity projection is intuitive for query-adaptive matching, it falls short in capturing local details. Decoupling the spatio-temporal modeling process in a divide-and-conquer manner proves more effective. Hence, we design progressive object-phrase prototype matching.
Constructing appearance prototype. For the given videos, we aggregate the patch features into object-level prototypes to represent fine-grained appearance representations, such as object instance, object part, and action region. During prototype construction, not all patch features are aggregated. Although the patch features contain important appearance representations, they also bring redundancy. For example, some background regions may interfere with cross-modal alignment. Hence, we filter out retrieval-superfluous information and generate object-level prototypes in a sparse aggregation manner. For convenience, three Fully Connected (FC) layers and ReLU function are utilized to predict sparse visual weights $W_{a}^{i}\in\mathbb{R}^{(C+1)\times N_{a}}$, where $N_{a}$ denotes the number of object-level prototypes in the $i$-th video. Therefore, we prevent these object-level prototypes from being affected by redundant patches.
For each frame $F_{v}^{i}\in\mathbb{R}^{(C+1)\times d}$, the constructed appearance prototype is defined as $P_{a}^{i}=W_{a}^{i^{T}}F_{v}^{i}\in\mathbb{R}^{N_{a}\times d}$. Ideally, each object prototype can adaptively aggregate the corresponding object-related or action-related patches. As for each phrase, we also apply a similar network structure to aggregate word features. Similarly, we utilize three FC layers and ReLU functions to obtain sparse textual weights $W_{p}^{j}\in\mathbb{R}^{(C+1)\times N_{p}}$, where $N_{p}$ is the number of phrase prototypes in the $j$-th query. Thus, we can fully extract the significant appearance representations by the fine-grained patch features and word features. Besides, the phrase prototypes $P_{p}^{i,j}=W_{p}^{j^{T}}F_{v}^{i}$ are optimized by spatial object-phrase prototype matching.
Object-phrase cross-modal alignment. We propose a prototype-wise query-video interaction from an appearance perspective. Specifically, we first compute the maximum similarity of object-phrase prototypes within each frame. This associates the phrase prototypes most similar to each object prototype, reflecting cross-modal fine-grained alignment. Then, for the multi-frame object similarity matrix, we find the largest similarity score across the frame sequences, which gives a more confident probability of object-phrase matching. Finally, the object-phrase matching scores are summed for the final similarity: $s_{a}=\frac{1}{N_{a}}\sum_{l=1}^{N_{a}}\max_{i=1}^{N_{v}}\max_{j=1}^{N_{p}}[P_{p}^{i,j}\times P_{a}^{l^{T}}]$.
Activity-Sentence Prototype Alignment
To model the motion information, we explore the activity-to-sentence relationship: 1) first perform progressive activity-sentence prototype aggregation to reveal the video semantic diversity, 2) then design dynamic prototype alignment.
Building activity prototype. To fully understand the video activity, we aim to design diverse activity prototypes. A naive solution to obtain video-level features based on global frame features is by mean pooling, or by adding motion encoder layers. However, this leads to two issues: 1) failure to perceive local details and ignoring important objects will exacerbate the bias of video feature learning; 2) these strategies generate a single video-level feature, which can only quantify one-to-one relations. Therefore, we investigate how to incorporate key fine-grained objects and dynamic motion changes into diverse activity prototypes.
The core idea is to progressively aggregate spatial object prototypes into frame prototypes and then perform inter-frame interaction to generate various activity prototypes. A frame decoder is first designed to incorporate all object prototypes $P_{a}\in\mathbb{R}^{(N_{v}\times N_{a})\times d}$ into frame-level prototypes $P_{f}\in\mathbb{R}^{N_{a}\times d}$, which implies fine-grained inter-object spatial relations. To learn frame-level object relations, we define the masked attention as: $P_{f}=Q_{f}+softmax(Q_{f}C_{a}^{T}+W_{f})V_{a}$, where $Q_{f}\in\mathbb{R}^{N_{a}\times d}$ refers to frame queries (i.e., a set of randomly initialized learnable features), $V_{a}$ and $C_{a}^{T}$ are the features after the linear transformation of object prototypes $P_{a}$. The attention mask $W_{f}\in\mathbb{R}^{N_{v}\times(N_{v}\times N_{a})}$ is:
$W_{f}(i,j)=\left\{\begin{matrix}0&\text{ if }N_{a}\cdot i\leq J<N_{a}\cdot(i+1),\\ -\infty&\text{otherwise.}\end{matrix}\right.$ (5)
We add frame prototype $p_{f}^{i}$ original global feature $v_{C}^{i}$ of corresponding frames to enhance the robustness of the model: $p_{f}^{i}=(p_{f}^{i}+v_{C}^{i})/{2}$.
Next, a dynamic activity decoder is developed to learn the inter-frame relationship in $P_{f}$, which can obtain different activity prototypes $P_{e}\in\mathbb{R}^{M_{q}\times d}$ to illustrate the rich information of videos. Our dynamic attention is formulated as: $P_{e}=Q_{e}+softmax(Q_{e}K_{a}^{T})V_{a}$, where $Q_{e}=[q_{e}^{1},q_{e}^{2},\cdots,q_{e}^{M_{q}}]\in\mathbb{R}^{M_{q}\times d}$ refers to activity queries, $V_{f}$ and $K_{f}^{T}$ are the features after the linear transformation of object prototypes $P_{f}$. During training, each activity query learns how to adaptively focus on video frame prototypes, while multiple queries implicitly guarantee a certain activity diversity. Differently, since the same video often corresponds to multiple text semantic descriptions, we directly use the global text representation $q_{e}$ as a sentence prototype to align with the activity prototypes $P_{e}$.
Activity-sentence alignment. The activity-sentence prototype alignment is expressed as: $s_{es}\!=\!\max_{i=1}^{M_{q}}(q_{e},P_{e_{i}})$. By the similarity, we can find the closest activity prototype to the text representation for dynamic alignment.
Multi-Modal Fusion and Grounding Head
[…]
In our model, we treat query-video pairs as positive examples, while considering all other pairwise combinations in the batch as negative examples. To fully leverage the query-video pair information, we introduce the query-to-video Robust InfoNCE (RINCE) loss as follows:
$\mathcal{L}_{q\rightarrow v}(S)\!=\!\frac{1}{M_{v}}\!\sum\nolimits_{i=1}^{M_{v}}[-\frac{\exp(S^{ii})}{\tau}\!+\!\frac{(\alpha\sum\nolimits_{j=1}^{M_{q}}\exp(S^{ij}))^{\tau}}{\tau}\!],$ (6)
where $\tau,\alpha\in(0,1]$ are learnable parameters. Similarly, the video-to-query loss is:
$\mathcal{L}_{v\rightarrow q}(S)\!=\!\frac{1}{M_{q}}\!\sum\nolimits_{j=1}^{M_{q}}[\!-\!\frac{\exp(S^{jj})}{\tau}\!+\!\frac{(\alpha\sum\nolimits_{i=1}^{M_{v}}\exp(S^{ji}))^{\tau}}{\tau}].$ (7)
Denoting activity-sentence and object-phrase prototype matching similarity matrices as $S_{es}$ and $S_{op}$ respectively, we design the alignment loss as follows:
$\mathcal{L}_{1}=\mathcal{L}_{v\rightarrow q}(S_{es})+\mathcal{L}_{v\rightarrow q}(S_{op})+\mathcal{L}_{q\rightarrow v}(S_{es})+\mathcal{L}_{q\rightarrow v}(S_{op}).$ (8)
To predict the segment start/end boundary quickly and accurately, we first introduce the span predictor (Zhang et al. 2020a) with two stacked transformer blocks and two feed-forward layers. Specifically, the multi-modal feature ${F}$ is fed into the span predictor, followed by a softmax function, to obtain two probability scores of start and end boundaries. We denote them as ${P}_{s(e)}\in\mathbb{R}^{N_{v}}$. The rounded integer boundaries $\hat{t}_{s(e)}$ are used to generate one-hot label vectors ${Y}_{s(e)}$ as supervision.
$\mathcal{L}_{2}=L_{CE}({P}_{s},{Y}_{s})+L_{CE}({P}_{e},{Y}_{e}),$ (9)
where $L_{CE}$ means the cross-entropy loss. The predicted boundary timestamps $\hat{t}_{s(e)}^{\prime}$ are obtained by $P_{s(e)}$: $(\hat{t_{s}}^{\prime},\hat{t_{e}}^{\prime})=arg\max_{\hat{t}_{s}^{\prime},\hat{t}_{e}^{\prime}}{P}_{s}(\hat{t_{s}}^{\prime}){P}_{e}(\hat{t}_{e}^{\prime})$, where $0\leq\hat{t}_{s}^{\prime}\leq\hat{t}_{e}^{\prime}\leq N_{v}$.
Since the above span predictor can only predict coarse integer boundary values, we additionally design a parallel float predictor consisting of several feed-forward layers to provide fine-grained float boundaries by the following loss:
$\mathcal{L}_{3}=f_{\text{L1-smooth}}(\hat{t}_{s}^{\prime},t_{s})+f_{\text{L1-smooth}}(\hat{t}_{e}^{\prime},t_{e}),$ (10)
where $f_{\text{L1-smooth}}$ represents the smooth L1 loss and $(t_{s},t_{e})$ is the ground-truth boundary. The predicted float boundaries ${O}_{s}$ and ${O}_{e}$ respectively represent the percentage of start and end boundary frames that are query-relevant. Therefore, the fine-grained boundary indexes $\tilde{t}_{s(e)}^{\prime}$ are calculated by: $(\tilde{t}_{s}^{\prime},\tilde{t}_{e}^{\prime})=(\hat{t}_{s}^{\prime}+1-{O}_{s},\hat{t}_{e}^{\prime}-1+{O}_{e})$.
The multi-modal network is trained by minimizing the weighted sum of the above losses, denoted as $\mathcal{L}$:
$\mathcal{L}=\mathcal{L}_{CL}+\lambda\mathcal{L}_{1}+\gamma\mathcal{L}_{2}+\mu\mathcal{L}_{3},$ (11)
where $\lambda,\gamma$ and $\mu$ are parameters to weigh different losses.
Experiments
ActivityNet Captions
Method Type R@1, R@1, R@5, R@5,
IoU=0.5 IoU=0.7 IoU=0.5 IoU=0.7
MRTNet FS 42.02 24.25 - -
2D-TAN FS 44.51 26.54 77.13 61.96
MMN FS 48.59 29.26 79.50 64.76
VDI FS 48.09 28.76 79.69 64.88
ICVC WS 29.52 - 66.61 -
VCA WS 31.00 - 53.83 -
WSTAN WS 30.01 - 63.42 -
CPL WS 31.37 - 43.13 -
MMDist WS 32.98 - - -
Ours MP 58.32 35.28 86.20 71.49
Charades-STA
Method Type R@1, R@1, R@5, R@5,
IoU=0.5 IoU=0.7 IoU=0.5 IoU=0.7
VDI FS 52.32 31.37 87.03 62.30
DRN FS 53.09 31.75 89.06 60.05
MESM FS 56.69 35.99 - -
MRTNet FS 62.50 43.63 - -
ICVC WS 31.02 16.53 77.53 41.91
VCA WS 38.13 19.57 78.75 37.75
LCNet WS 39.19 18.17 80.56 45.24
CPL WS 49.24 22.39 84.71 52.37
MMDist WS 54.72 26.00 - -
Ours MP 69.93 46.27 97.16 69.13
TACoS
Method Type R@1, R@1, R@5, R@5,
IoU=0.3 IoU=0.5 IoU=0.3 IoU=0.5
DRN FS - 23.17 - 33.36
2D-TAN FS 37.29 25.32 57.81 45.04
MRTNet FS 37.81 26.01 - -
MMN FS 39.24 26.17 62.03 47.39
MIGCN FS 48.79 37.57 67.63 57.91
MESM FS 52.69 39.52 - -
Ours MP 53.38 42.62 73.54 62.24
Table 1: Effectiveness comparison for TSG on all the datasets under official train/test splits, where “MP” means “Multi-Pair”.
Datasets. For a fair comparison with previous works (Zhang et al. 2020b; Wang et al. 2022c; Zheng et al. 2022b), we utilize the same ActivityNet Captions (Caba Heilbron et al. 2015), TACoS (Regneri et al. 2013), and Charades-STA (Sigurdsson et al. 2016) datasets for evaluation.
Evaluation metrics. Following (Gao et al. 2017), we evaluate the grounding performance by “R@n, IoU=m”, which means the percentage of queries having at least one result whose Intersection over Union (IoU) with ground truth is larger than m.
Compared methods. 1) Fully-supervised (FS): DRN (Zeng et al. 2020), 2D-TAN (Zhang et al. 2020b), MIGCN (Zhang et al. 2021), MMN (Wang et al. 2022c), VDI (Luo et al. 2023), MRTNet (Ji et al. 2024c), MESM (Liu et al. 2024e). 2) Weakly-supervised (WS): LCNet (Yang et al. 2021), VCA (Wang, Chen, and Jiang 2021), ICVC (Chen et al. 2022), WSTAN (Wang et al. 2022b), CNM (Zheng et al. 2022a), CPL (Zheng et al. 2022b), MMDist (Bao et al. 2024).
Methods Variant R@1, R@1, R@5, R@5, VPS
IoU=0.5 IoU=0.7 IoU=0.5 IoU=0.7
2D-TAN Origin 39.81 23.25 79.33 52.15 12.43
+Ours 43.37 27.48 86.74 55.68 46.28
MMN Origin 47.31 27.28 83.74 58.41 21.63
+Ours 49.36 31.03 88.75 62.39 48.38
LCNet Origin 39.19 18.17 80.56 45.24 23.88
+Ours 42.80 25.94 82.46 51.37 52.73
Table 2: We serve our method as a plug-and-play module for state-of-the-art TSG methods on Charades-STA under official train/test splits, where “VPS” denotes “video per second” during inference.
Figure 3: Performance comparison with state-of-the-art methods on Charades-STA. (a) compares the effectiveness (R@1, IoU=0.5) and the efficiency (QPS), (b) shows that our method can serve as a plug-and-play module to enhance their efficiency, (c) is the qualitative results.
Model ActivityNet Captions Charades-STA TACoS
R@1 R@1 R@5 R@5 R@1 R@1 R@5 R@5 R@1 R@1 R@5 R@5
IoU=0.5 IoU=0.7 IoU=0.5 IoU=0.7 IoU=0.5 IoU=0.7 IoU=0.5 IoU=0.7 IoU=0.3 IoU=0.5 IoU=0.3 IoU=0.5
w/o CSM 53.82 28.60 81.95 67.76 65.82 44.45 92.11 63.19 47.38 33.12 68.44 56.13
w/o OPM 55.40 30.33 83.95 70.02 59.80 39.13 93.32 62.01 49.85 37.24 69.48 59.16
w/o APA 56.27 30.86 80.21 68.49 60.77 42.96 90.11 65.24 50.92 37.25 70.42 59.08
w/o AVM 54.95 31.03 83.72 69.30 65.21 45.75 93.24 67.69 50.46 38.05 70.85 58.49
Full 58.32 35.28 86.20 71.49 69.93 46.27 97.16 69.13 53.38 42.62 73.54 62.24
Table 3: Main ablation study, where we remove each key individual component to investigate its effectiveness. “CSM” denotes “cross-sentence semantic mining”, “OPM” denotes “object-phrase prototype matching”, “APA” denotes “activity-sentence prototype alignment”, “AVM” denotes “adaptive video-query matching”, “Full” denotes our full model.
Comparison With State-Of-The-Arts
Quantitative comparison. We compare our proposed method with other existing state-of-the-art approaches in Table 1 and Figure 3(a). Obviously, our proposed method outperforms both fully-supervised and weakly-supervised methods by a large margin. The main reasons are as follows: 1) ActivityNet Captions: these sentence annotations share many nouns and sequencing words, and our method can mine the spatial and temporal relationship between different sentences by our cross-sentence semantic mining module. 2) Charades-STA: some video-query pairs share the same sentence queries. Thus, our method can co-train these pairs and transfer the knowledge from some easy pairs to difficult pairs. 3) TACoS: TACoS only contains activities of cooking scenarios, where these videos often share similar object/appearance information. Our method leverages the information to co-train different videos for better video understanding.
Plug-and-play. Besides, our method can serve as a plug-and-play module for state-of-the-art models (fully-supervised TSG: 2D-TAN and MMN, weakly-supervised TSG: LCNet). As shown in Table 2 and Figure 3(b), our method can significantly improve their performance with higher efficiency on the Charades-STA dataset. Therefore, our multi-thread framework for TSG is flexible and can be adopted into other state-of-the-art methods to improve their effectiveness and efficiency. The performance improvement is because 1) effectiveness: our multi-thread framework can mine the semantic relationship between different queries, which allows them to assist each other for grounding. 2) efficiency: our framework can co-train multiple video-query pairs to reduce redundant calculations.
Visualization comparison. To qualitatively investigate the effectiveness of our method, we report some representative examples in Figure 3(c), where our grounding result is closer to the ground truth than MESM and WSTAN.
Ablation Study
Main ablation studies. To evaluate the contribution of each module, we perform the main ablation study in Table 3. All the modules contribute a lot to the final performances, demonstrating their effectiveness in exploring the intra- and inter-modal relationship in multiple video-query pairs.
Model R@1 R@1 R@5 R@5
IoU=0.5 IoU=0.7 IoU=0.5 IoU=0.7
w/o Temporal relationship 56.40 30.98 78.36 69.88
w/o Contextual relationship 57.14 31.06 80.12 67.84
w/o Semantic relationship 54.85 30.81 78.48 68.95
Full 58.32 35.28 86.20 71.49
Table 4: Ablation study on cross-sentence semantic mining.
Importance of cross-sentence semantic mining (CSM). To assess the effectiveness of our CSM module, we compare different ablation models on ActivityNet Captions in Table 4, where we remove one relationship between different sentence queries in the first three ablation models. Obviously, our full model achieves the best performance because ActivityNet Captions contains a large number of semantically related queries, and our model can fully mine the semantic relationship between different queries for grounding.
Cross-modal contrast via self-supervision
Visual Textual R@1 R@1 R@5 R@5
feature feature IoU=0.5 IoU=0.7 IoU=0.5 IoU=0.7
53.03 29.96 80.31 68.69
52.84 30.12 79.68 67.45
58.32 35.28 86.20 71.49
Adaptive negative selection
Fixed threshold 54.08 31.09 82.20 65.13
w/o $\cos(p\pi+1)$ 55.15 31.26 82.39 68.04
w/ $\cos(p\pi+1)$ 58.32 35.28 86.20 71.49
Table 5: Ablation study on adaptive video-query matching.
Influence of cross-modal contrast (CC). To analyze the importance of our CC module, we conduct corresponding experiments on ActivityNet Captions in Table 5. Both visual and textual features contribute a lot to integrating different video-query pairs. It is because our model can use the self-supervision strategy in Eq. (Adaptive Video-Query Matching) to fully mine the intra- and inter-modal representations for video grounding.
Effect of adaptive negative selection (ANS). To assess the performance of our ANS module, we change the threshold to obtain two ablation models in Table 5. Obviously, our full model obtains the best results since our ANS module can generate an adaptive threshold for negative query selection, which can fully match queries and relevant videos.
CAP OCA R@1 R@1 R@5 R@5
IoU=0.5 IoU=0.7 IoU=0.5 IoU=0.7
56.80 31.03 83.16 70.93
57.03 30.72 84.13 69.48
58.32 35.28 86.20 71.49
Table 6: Ablation study on object-phrase prototype matching, where “CAP” denotes “constructing appearance prototype” and “OCA” denotes “object-phrase cross-modal alignment”.
Significance of object-phrase prototype matching (OPM). To analyze the performance of our OPM module in integrating the cross-modal spatial information, an ablation experiment is conducted on ActivityNet Captions in Table 6. Our full model beats other ablation models by a large margin since appearance and phrase prototypes provide visual and textual spatial semantics for multi-modal fusion.
BMP AA R@1 R@1 R@5 R@5
IoU=0.5 IoU=0.7 IoU=0.5 IoU=0.7
54.88 30.91 82.42 67.60
53.04 31.09 83.86 68.95
58.32 35.28 86.20 71.49
Table 7: Ablation study on activity-sentence prototype alignment, where “BMP” denotes “building motion prototype” and “AA” denotes “activity-sentence alignment”.
Analysis on activity-sentence prototype alignment (APA). We further analyze the performance of our APA module for cross-modal temporal representations on ActivityNet Captions in Table 7. Obviously, both modules bring significant performance improvement since activity and sentence prototypes can be used to understand the temporal semantics from visual and textual perspective respectively.
Conclusion
In this paper, we pose a brand-new and realistic setting: Multi-Pair TSG. For the challenging task, we propose a novel Multi-Thread Knowledge Transfer Network (MKTN) to deeply explore intra- and inter-modal relationships. Extensive experiments on three challenging datasets show the effectiveness and efficiency of our MKTN. Moreover, our MKTN can serve as a plug-and-play module for previous methods to enhance their effectiveness and efficiency.